graph the equation of this circle
1/4x^2+1/4y^2=16

Answer by ewatrrr(24785)

 
Hi,
1/4x^2+1/4y^2=16
0r
x^2 + y^2 = 64
x^2 + y^2 = 8^2 C(0,0) r = 8